# Large Language Model Agent for Modular Task Execution in Drug Discovery

### v1

@@ 4 Discussion
For molecule generation, AgentD integrates external models such as REINVENT and Mol2Mol using natural language–driven configuration updates, enabling seamless integration of new tools with minimal customization. By leveraging a language model to interpret and modify configuration files based on high-level instructions, the system can flexibly switch between different generative models or sampling strategies without manual code changes. This approach removes the need for model-specific hardcoded logic and allows for rapid adaptation to evolving architectures and file structures. The modular design of AgentD further ensures its long-term adaptability as generative modeling techniques continue to advance.
The property-aware molecular refinement module shows both promise and limitations. AgentD successfully improves target properties such as permeability or toxicity using only its built-in reasoning and domain knowledge. However, molecular optimization remains intrinsically complex due to interdependencies among properties. For example, in Figure 6b, improvements in permeability (logPapp) lead to declines in lipophilicity (logP) and solubility (logS). Similarly, the first-round refinement in Figure 6c mitigates predicted liver toxicity but worsens permeability. These trade-offs highlight the fundamental challenge of multi-property optimization in drug design. Despite these complexities, AgentD achieves a net improvement in overall drug-likeness across the molecule pool, as shown in Figure 7c and d. These findings highlight the importance of incorporating multi-objective optimization in future iterations to better balance competing pharmacological goals.
The structure generation task complements the pooling and refinement process by enabling 3D protein-ligand complex generation, which provides critical input for structure-based studies. Although we use general drug-likeness criteria such as Lipinski’s Rule of Five and QED thresholds for demonstration purposes, target-specific filters can be readily implemented in practice. Boltz-2 captures meaningful trends in ligand binding behavior, but its IC50 and inhibitor probability predictions should be interpreted as rough approximations that are useful for initial prioritization rather than as substitutes for detailed structure-based evaluation.
@@ 6.1 Large Language Model Agent
Our framework employs OpenAI’s GPT-4o model 51 as the primary language model for all agentD modules. GPT-4o is an optimized version of GPT-4, which belongs to the Generative Pretrained Transformer family 52. The model uses transformer architecture with self-attention mechanisms 53 to effectively process contextual relationships in text sequences. GPT-4o maintains the strong language understanding capabilities of GPT-4 while offering improved computational efficiency through performance optimizations. We selected GPT-4o specifically for its proven effectiveness in scientific reasoning tasks and its extensive training on chemical and biological literature. We implement the agent framework using LangChain 54, a Python library for developing large language model applications. LangChain simplifies the management of complex workflows by handling prompt coordination, response processing, and integration with external tools and APIs.
@@ 6.2 Database
To facilitate target-specific molecular design, we utilize two established bioinformatics resources, UniProt and ChEMBL, to access protein sequence data and small molecule representations, respectively. UniProt provides high-quality annotated protein sequences between species, enabling access to functionally characterized targets of therapeutic relevance 34. For each therapeutically relevant protein target (e.g., EGFR or TP53), the agent queries UniProt’s REST API, restricted to the Homo sapiens taxonomy (organism ID: 9606) to retrieve the corresponding amino acid sequence in FASTA format. When existing drugs are available for the target, we subsequently query ChEMBL to obtain SMILES representations of these known compounds35. This ChEMBL query step is only executed if existing drugs for the target are identified through prior web searches; otherwise, this step is skipped.
@@ 6.3 ADMET Properties
ADMET properties encompass Absorption, Distribution, Metabolism, Excretion, and Toxicity characteristics that are critical for drug development 55, 56. These pharmacologically relevant descriptors determine how a compound is absorbed into the bloodstream, distributed across tissues, metabolized by enzymatic systems, eliminated from the body, and whether it poses potential toxic effects. Early prediction of ADMET properties is essential for prioritizing compounds with favorable biopharmaceutical and safety profiles, thereby reducing downstream attrition during drug development 57, 58.
As part of this multi-stage assessment, we integrated Deep-PK, an ADMET prediction framework that operates on SMILES input and internally employs a Message Passing Neural Network (MPNN) to capture the atomic and topological features of each molecule 39. This architecture allows for graph-based encoding of chemical structures by propagating information across atom–bond interactions. Candidate ligands, whether recovered from ChEMBL or generated, are submitted to the Deep-PK REST API via their SMILES strings, and the resulting ADMET profiles are parsed to prioritize compounds with favorable pharmacokinetic and safety attributes. The complete set of predicted properties was summarized in Supplementary Information Table S2.
@@ 6.4 Binding Affinity
Binding affinity to the biological target is a fundamental determinant of therapeutic potential, serving as a key predictor of drug potency and selectivity. We employ two complementary strategies for affinity prediction. In the property prediction task, the sequence-based BAPULM model employs a dual encoder architecture to estimate the dissociation constant ($K_{d}$) from protein amino acid sequences and ligand SMILES representations. It integrates two domain-specific pre-trained language models: ProtT5-XL-U50 for proteins and MolFormer for small molecules 59, 60. Each encoder generates latent embeddings tailored to its respective input, which are then projected into a shared latent space using learnable feedforward projection heads. Subsequently, a predictive head processes these joint representations to estimate the binding affinity, reported as $pK_{d}=-\log_{10}(K_{d})$40. Given its computational efficiency and reliance solely on sequence-level inputs, it is well-suited for early-stage screening of large ligand libraries based on predicted binding affinity.
In the structure generation task, the structure-based model Boltz begins with the same sequence and SMILES input but internally generates 3D protein–ligand complex structures. These conformations are then used to predict the half-maximum inhibitory concentration (IC50), reflecting the inhibitory efficacy of a compound in biochemical assays 61. In addition to regression-based affinity values, Boltz also outputs inhibitor probability scores, indicating the likelihood that a given ligand acts as an active binder. Although $K_{d}$ and IC50 originate from different experimental setups, they are related through the Cheng-Prusoff equation 62.
@@ 6.5 SMILES Generation
We utilize two molecular generators: REINVENT for de novo molecular design and Mol2Mol for molecular optimization 37, 36, 38. Both generators utilize recurrent neural networks and transformer architectures and are embedded within machine learning optimization algorithms, including reinforcement learning and transfer learning 37, 63.
REINVENT performs de novo molecular generation using sequence-based models that capture SMILES token probabilities in an autoregressive manner. The models are trained via teacher-forcing on large SMILES datasets to learn chemical syntax and generate valid molecules 64. Reinforcement learning optimization employs a policy gradient scheme using the “Difference between Augmented and Posterior” (DAP) strategy, where scalar scores are computed for generated molecules based on user-defined scoring functions 65. An augmented likelihood combines this reward signal with a fixed prior model likelihood, providing regularization to maintain chemical plausibility while optimizing toward desired properties.
Mol2Mol performs conditional generation, accepting input SMILES strings and generating structurally similar molecules within user-defined similarity constraints 38. The transformer-based model was trained on over 200 billion molecular pairs from PubChem with Tanimoto similarity $\geq$ 0.50. Training employed ranking loss to directly link negative log-likelihood to molecular similarity. The model supports multinomial sampling with temperature control and beam search decoding, enabling controlled exploration around known compounds while preserving structural relationships to the input molecule.
@@ 6.6 Protein-Ligand Complex Structure
Boltz-2 is a structure-based deep learning model designed to jointly predict 3D protein–ligand complex structures by integrating protein folding and ligand binding into a unified framework 42. The model takes as input a protein FASTA sequence and a ligand SMILES string, and simultaneously infers the full atomic conformation of the protein as well as the bound pose of the ligand within the predicted binding pocket. Unlike traditional docking pipelines that require experimentally resolved protein structures, Boltz-2 performs ab initio structure prediction, enabling end-to-end modeling from sequence alone.
Its architecture builds upon the Evoformer17 stack and SE(3)-equivariant transformer modules, incorporating interleaved attention mechanisms to capture long-range dependencies both within the protein sequence and between the protein and ligand 53, 41, 42. This allows Boltz-2 to reason over flexible ligand conformations and protein-side chain rearrangements in a physics-aware manner. By directly predicting all-atom 3D coordinates, the model enables rapid generation of realistic protein–ligand complexes suitable for downstream scoring.
@@ 6.7 Evaluation Metrics
To evaluate the drug-likeness and pharmacokinetic relevance of the generated molecules, we applied five widely used rule-based filters: Lipinski’s Rule of Five 46, Veber’s Rule 47, Ghose Filter 48, Rule of Three (Ro3)49, and Oprea’s Lead-like Rule50. The specific numerical criteria for each rule are summarized in Table S3 of the Supporting Information. Molecular descriptors were computed using RDKit where necessary.
Lipinski’s Rule of Five includes five criteria: molecular weight (MW) $\leq$ 500 Da, LogP $\leq$ 5, hydrogen bond donors (HBD) $\leq$ 5, hydrogen bond acceptors (HBA) $\leq$ 10, and rotatable bonds $\leq$ 10. Although the original rule included only four properties, the rotatable bond constraint is now widely adopted to better capture oral bioavailability. A molecule was considered compliant if it satisfied at least four of the five conditions.
@@
The Ghose filter identifies well-balanced drug-like molecules based on MW between 160–480 Da, LogP between –0.4 and 5.6, molar refractivity (MR) between 40–130, and heavy atom count between 20–70.
The Rule of Three, designed for fragment-based discovery, applies stricter limits: MW $<$ 300 Da, LogP $\leq$ 3, and HBD $\leq$ 3. It is used to identify small, synthetically accessible fragments with potential for optimization.
Oprea’s Lead-like Rule defines criteria suitable for early-stage lead optimization: MW between 200–450 Da, LogP between –1 and 4.5, HBD $\leq$ 5, HBA $\leq$ 8, rotatable bonds $\leq$ 8, and aromatic rings $\leq$ 4. One violation among these conditions was permitted to allow flexibility in lead selection.

### v2

@@ 4 Discussion
For molecule generation, AgentD integrates external models such as REINVENT and Mol2Mol using natural language-driven configuration updates, enabling seamless integration of new tools with minimal customization. By leveraging a language model to interpret and modify configuration files based on high-level instructions, the system can flexibly switch between different generative models or sampling strategies without manual code changes. This approach removes the need for model-specific hardcoded logic and allows for rapid adaptation to evolving architectures and file structures. The modular design of AgentD further ensures its long-term adaptability as generative modeling techniques continue to advance.
The property-aware molecular refinement module shows both promise and limitations. AgentD successfully improves target properties such as permeability or toxicity using only its built-in reasoning and domain knowledge. However, molecular optimization remains intrinsically complex due to interdependencies among properties. For example, in Figure 6a and b, predicted apparent permeability (logPapp) is influenced by lipophilicity (logP), consistent with Overton’s rule 61, whereas aqueous solubility (logS) does not follow a consistent trend and may either increase or decrease. These observations highlight the complex, non-linear interplay among lipophilicity, solubility, and permeability. Similarly, the first-round refinement in Figure 6d mitigates predicted liver toxicity but worsens permeability. These trade-offs highlight the fundamental challenge of multi-property optimization in drug design. These findings highlight the importance of incorporating multi-objective optimization in future iterations to better balance competing pharmacological goals.
Despite these complexities, AgentD achieves a net improvement in overall drug-likeness across the molecule pool, as shown in Figure 7c and d. Notably, the most significant improvement is observed for the Ghose filter. Because the Ghose filter reflects physicochemical property ranges derived from known drugs, this trend suggests that the LLM may leverage its training knowledge of existing molecules when refining candidates. In contrast, performance on Lipinski’s Rule of Five slightly deteriorates. Since this rule emphasizes oral bioavailability, and oral absorption was rarely designated as a primary weakness in the molecular refinement tasks (Table S3), this limited improvement is not unexpected.
The structure generation task complements the pooling and refinement process by enabling 3D protein-ligand complex generation, which provides critical input for structure-based studies. Boltz-2 captures meaningful trends in ligand binding behavior, but its IC50 and inhibitor probability predictions should be interpreted as rough approximations that are useful for initial prioritization rather than as substitutes for detailed structure-based evaluation.
@@ 6.1 Large Language Model Agent
Our framework employs OpenAI’s GPT-4o model 62 as the primary language model for all agentD modules. GPT-4o is an optimized version of GPT-4, which belongs to the Generative Pretrained Transformer family 63. The model uses transformer architecture with self-attention mechanisms 64 to effectively process contextual relationships in text sequences. GPT-4o maintains the strong language understanding capabilities of GPT-4 while offering improved computational efficiency through performance optimizations. We selected GPT-4o specifically for its proven effectiveness in scientific reasoning tasks and its extensive training on chemical and biological literature. We implement the agent framework using LangChain 65, a Python library for developing large language model applications. LangChain simplifies the management of complex workflows by handling prompt coordination, response processing, and integration with external tools and APIs. Accessing the LLM API does not require local GPU resources. For tool-specific tasks such as binding affinity prediction, SMILES generation, and protein–ligand structure generation, we utilized a single NVIDIA RTX 2080 Ti GPU.
@@ 6.2 Database
To facilitate target-specific molecular design, we utilize two established bioinformatics resources, UniProt and ChEMBL, to access protein sequence data and small molecule representations, respectively. UniProt provides high-quality annotated protein sequences between species, enabling access to functionally characterized targets of therapeutic relevance 42. For each therapeutically relevant protein target (e.g., EGFR or TP53), the agent queries UniProt’s REST API, restricted to the Homo sapiens taxonomy (organism ID: 9606) to retrieve the corresponding amino acid sequence in FASTA format. When existing drugs are available for the target, we subsequently query ChEMBL to obtain SMILES representations of these known compounds43. This ChEMBL query step is only executed if existing drugs for the target are identified through prior web searches; otherwise, this step is skipped.
@@ 6.3 ADMET Properties
ADMET properties encompass Absorption, Distribution, Metabolism, Excretion, and Toxicity characteristics that are critical for drug development 66, 67. These pharmacologically relevant descriptors determine how a compound is absorbed into the bloodstream, distributed across tissues, metabolized by enzymatic systems, eliminated from the body, and whether it poses potential toxic effects. Early prediction of ADMET properties is essential for prioritizing compounds with favorable biopharmaceutical and safety profiles, thereby reducing downstream attrition during drug development 68, 69.
As part of this multi-stage assessment, we integrated Deep-PK, an ADMET prediction framework that operates on SMILES input and internally employs a Message Passing Neural Network (MPNN) to capture the atomic and topological features of each molecule 47. This architecture allows for graph-based encoding of chemical structures by propagating information across atom–bond interactions. Candidate ligands, whether recovered from ChEMBL or generated, are submitted to the Deep-PK REST API via their SMILES strings, and the resulting ADMET profiles are parsed to prioritize compounds with favorable pharmacokinetic and safety attributes. The complete set of predicted properties is summarized in Supplementary Information Table S2.
@@ 6.4 Binding Affinity
Binding affinity to the biological target is a fundamental determinant of therapeutic potential, serving as a key predictor of drug potency and selectivity. We employ two complementary strategies for affinity prediction. In the property prediction task, the sequence-based BAPULM model employs a dual encoder architecture to estimate the dissociation constant ($K_{d}$) from protein amino acid sequences and ligand SMILES representations. It integrates two domain-specific pre-trained language models: ProtT5-XL-U50 for proteins and MolFormer for small molecules 70, 71. Each encoder generates latent embeddings tailored to its respective input, which are then projected into a shared latent space using learnable feedforward projection heads. Subsequently, a predictive head processes these joint representations to estimate the binding affinity, reported as $pK_{d}=-\log_{10}(K_{d})$48. Given its computational efficiency and reliance solely on sequence-level inputs, it is well-suited for early-stage screening of large ligand libraries based on predicted binding affinity.
In the structure generation task, the structure-based model Boltz begins with the same sequence and SMILES input but internally generates 3D protein–ligand complex structures. These conformations are then used to predict the half-maximum inhibitory concentration (IC50), reflecting the inhibitory efficacy of a compound in biochemical assays 72. In addition to regression-based affinity values, Boltz also outputs inhibitor probability scores, indicating the likelihood that a given ligand acts as an active binder. Although $K_{d}$ and IC50 originate from different experimental setups, they are related through the Cheng-Prusoff equation 73.
@@ 6.5 SMILES Generation
We utilize two molecular generators: REINVENT for de novo molecular design and Mol2Mol for molecular optimization 45, 44, 46. Both generators utilize recurrent neural networks and transformer architectures and are embedded within machine learning optimization algorithms, including reinforcement learning and transfer learning 45, 74.
REINVENT performs de novo molecular generation using sequence-based models that capture SMILES token probabilities in an autoregressive manner. The models are trained via teacher-forcing on large SMILES datasets to learn chemical syntax and generate valid molecules 75. Reinforcement learning optimization employs a policy gradient scheme using the “Difference between Augmented and Posterior” (DAP) strategy, where scalar scores are computed for generated molecules based on user-defined scoring functions 76. An augmented likelihood combines this reward signal with a fixed prior model likelihood, providing regularization to maintain chemical plausibility while optimizing toward desired properties.
Mol2Mol performs conditional generation, accepting input SMILES strings and generating structurally similar molecules within user-defined similarity constraints 46. The transformer-based model was trained on over 200 billion molecular pairs from PubChem with Tanimoto similarity $\geq$ 0.50. Training employed ranking loss to directly link negative log-likelihood to molecular similarity. The model supports multinomial sampling with temperature control and beam search decoding, enabling controlled exploration around known compounds while preserving structural relationships to the input molecule.
@@ 6.6 Protein-Ligand Complex Structure
Boltz-2 is a structure-based deep learning model designed to jointly predict 3D protein–ligand complex structures by integrating protein folding and ligand binding into a unified framework 50. The model takes as input a protein FASTA sequence and a ligand SMILES string, and simultaneously infers the full atomic conformation of the protein as well as the bound pose of the ligand within the predicted binding pocket. Unlike traditional docking pipelines that require experimentally resolved protein structures, Boltz-2 performs ab initio structure prediction, enabling end-to-end modeling from sequence alone.
Its architecture builds upon the Evoformer19 stack and SE(3)-equivariant transformer modules, incorporating interleaved attention mechanisms to capture long-range dependencies both within the protein sequence and between the protein and ligand 64, 49, 50. This allows Boltz-2 to reason over flexible ligand conformations and protein-side chain rearrangements in a physics-aware manner. By directly predicting all-atom 3D coordinates, the model enables rapid generation of realistic protein–ligand complexes suitable for downstream scoring.
@@ 6.7 Evaluation Metrics
To evaluate the drug-likeness, lead-likeness, and pharmacokinetic relevance of the generated molecules, we applied four widely used rule-based filters: Lipinski’s Rule of Five 57, Veber’s Rule 58, the Ghose filter 59, and Oprea’s Lead-like Rule 60. The first three rules assess general drug-likeness, while Oprea’s rule specifically addresses lead-likeness as an early-stage starting point for optimization. The detailed numerical thresholds for each rule are summarized in Table S4 of the Supporting Information. All descriptors were computed using RDKit.
Lipinski’s Rule of Five includes five criteria: molecular weight (MW) $\leq$ 500 Da, LogP $\leq$ 5, hydrogen bond donors (HBD) $\leq$ 5, hydrogen bond acceptors (HBA) $\leq$ 10, and rotatable bonds $\leq$ 10. A molecule was considered compliant if it satisfied at least four of the five conditions. Although the original formulation included only the first four parameters, the rotatable bond threshold has since been adopted to better capture oral absorption potential.
@@
The Ghose filter identifies drug-like chemical space based on statistical distributions observed in marketed drugs. Its thresholds include MW between 160–480 Da, LogP between –0.4 and 5.6, molar refractivity (MR) between 40–130, and heavy atom count between 20–70. Molecules meeting these ranges are considered to occupy favorable physicochemical space for drug development.
Oprea’s Lead-like Rule defines chemical properties suitable for “lead” compounds—molecules that may not yet meet full drug-likeness but serve as starting points for further optimization. Criteria include MW between 200–450 Da, LogP between –1 and 4.5, HBD $\leq$ 5, HBA $\leq$ 8, rotatable bonds $\leq$ 8, and aromatic rings $\leq$ 4, with allowance for one violation. Compared to drug-like rules, these thresholds are narrower, reflecting the preference for smaller, less lipophilic molecules that can be synthetically elaborated into drug candidates.